OpenStudy (anonymous):

find a formula for the inverse g^-1(x) of the function g(x)=x^3+1

OpenStudy (anonymous):

y= x^3 +1 x^3=y-1 x=cuberoot(y-1) g^-1(x)= cuberoot(x-1)

OpenStudy (anonymous):

G^-1 X =cube root (x-1)